Grade 10General Physics

when going east at 10 kmph a train moving with const. velocity appears to be moving exactly north east . when my velocity is increased to 30 kmph east it appears to be moving north. a.what is the velocity of train in noth? b.what is the velocity of train along east? ​c.what speed should i move north so that train appears to move exactly south-east? please provide commplete solution in details.. :)

Akshit Bansal
9 Years agoGrade 10

1 Answer

Vikas TU
9 Years ago
We can solve it by relative speed: 
Speed of man = 10i 
Speed of prepare = xi+yj 
Relative speed = (x-10)i+yj 
Since both part to be equivalent to be in north east 
x-10 = y 
Presently 
Relative vel = (x-30)i+yj 
    X =30 y =20 
Speed of prepare = 30i+20j 
= 10√13